# Convolutional L2LFlows: Generating Accurate Showers in Highly Granular Calorimeters Using Convolutional Normalizing Flows

### v2

@@ 1 Introduction
However, the connections per calorimeter layer remained relatively naive. In this work, we investigate two key innovations to improve the scalability and fidelity of normalizing flows: i) convolutional layers assuming a approximate translational symmetry in lateral direction; ii) skip connections inspired by the U-net architecture [110] to further improve the flow of information and network convergence; and iii) coupling layers instead of masked autoregressive flows for fast training and inference [111, 112, 113, 114, 35] without the need of probability density distillation.
@@ 3.1 Energy Distribution Flow
$p(E_{1},\dots,E_{N}|E_{inc})$ (2)
where $E_{1}$ to $E_{N}$ denote the layer energies and $E_{inc}$ denotes the incident energy. As this energy distribution flow is not the bottleneck of the generative task, we use the same preprocessing and architecture as in [31]. It consists of 6 MADE blocks [127, 128] with rational quadratic splines (RQS) [114]. We apply permutation layers between these MADE blocks. A logit transformation is used to preprocess the layer energies, and a log transformation is used for the incident energy.
@@ 3.3 Training
Calorimeter data are sparse, meaning that most input values are zero. This poses a problem for training flows. Since the data points live in a submanifold of the data space, the density and therefore the loss are not bounded. This problem is usually solved by adding noise. As in the original CaloFlow, we add uniform noise between zero and $1$ keV to all voxels. In addition, we fill zero voxels with log Gaussian distributed values to avoid sharp edges.
We trained one instance of the networks for 200 epochs on GettingHigh and one instance for 800 epochs on CaloChallenge dataset 3. The different amount of epochs is necessary because of the different training statistics. We applied a one-cycle learning rate scheduler [133], decoupled L2 regularization [134], and L2 gradient clipping. ADAM [131] was used as optimizer. As data augmentation, we rotated the CaloChallenge dataset 3 showers by random angles in the $x-y$ plane. This is possible because the incident particles enter the calorimeter perpendicular to its surface and no magnetic field brakes the symmetry. Further details on architecture and training are provided in A.
@@ 3.4 Postprocessing
It is not guaranteed that the causal flows produces showers that match the layer energies generated by the energy distribution flow exactly. For that reason,[27, 31] rescaled showers per-layer to match the target. However, like the authors of [33], we observe that rescaling generated CaloChallenge showers results in an excess of low energy hits due to zero voxels getting shifted above the threshold. Therefore, instead of applying postprocessing for CaloChallenge, the showers generated by the causal flows are used directly. For GettingHigh we use the same energy rescaling as in [31]. See Appendix B.3 in [31] for a detailed description.
@@ 4.2 Classifier
Histograms and one-dimensional Wasserstein distances are useful to test how well one-dimensional distributions match, but they do not account for correlations. Instead, we use classifier tests to compare generative models based on higher-dimensional distributions [27, 28, 33, 31].
@@ 5.1 Observables
The diagrams in Figure 9 show numerical evaluated one-dimensional Wasserstein distances between test data and generated data. Given are the Wasserstein distances for all observables described in section 4.1. To make the Wasserstein distance dimensionless, the distributions are standard scaled using the mean and standard deviation of the Geant4 distribution. We see a clear improvement compared to the BIB-AE in almost all observables. We also notice that for many observables, statistics are insufficient to resolve any differences between Geant4 and Flow-generated samples.
@@ Acknowledgments
This research was supported in part by the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany. This project has received funding from the European Union’s Horizon 2020 Research and Innovation programme under Grant Agreement No 101004761. We acknowledge support by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and via the KISS consortium (05D23GU4, 13D22CH5) funded by the German Federal Ministry of Education and Research BMBF in the ErUM-Data action plan. CK would like to thank the Baden-Württemberg-Stiftung for financing through the program Internationale Spitzenforschung, project Uncertainties – Teaching AI its Limits (BWST_IF2020-010).
@@ Appendix B Preprocessing
$\text{logit}_{\alpha}(x)$ $=\log\left(\frac{x\times(1-2\alpha)+\alpha}{1-x\times(1-2\alpha)-\alpha}\right)$ (5)
@@
$E_{\text{max}}$ $=\max_{\text{training data}}\max_{i}E_{i}$ (6)
$x_{i}$ $=\text{logit}_{10^{-6}}\left(\frac{E_{i}}{E_{\text{max}}}\right)$ (7)
@@
$I_{\text{max},i}$ $=\max_{\text{training data}}\max_{jk}I_{ijk}$ (8)
$x_{ijk}$ $=\text{logit}_{10^{-2}}\left(1+\frac{\log\left(10^{-7}+\frac{I_{ijk}}{v_{\text{max},i}}\right)}{\log(10^{7})}\right)$ (9)
@@
$\text{condition}_{inc}$ $=\log\left(10^{-6}+\frac{E_{inc}}{1~{}\text{GeV}}\right)$ (10)
$\text{condition}_{i}$ $=\log\left(10^{-6}+\frac{E_{i}}{1~{}\text{GeV}}\right)$ (11)

### v3

@@ 1 Introduction
However, the connections per calorimeter layer remained relatively naive. In this work, we investigate three key innovations to improve the scalability and fidelity of normalizing flows: i) convolutional layers assuming a approximate translational symmetry in lateral direction; ii) skip connections inspired by the U-net architecture [110] to further improve the flow of information and network convergence; and iii) coupling layers instead of masked autoregressive flows for fast training and inference [111, 112, 113, 114, 35] without the need of probability density distillation.
@@ 3.1 Energy Distribution Flow
$p(E_{1},\dots,E_{N}|E_{inc})$ (3.2)
where $E_{1}$ to $E_{N}$ denote the layer energies and $E_{inc}$ denotes the incident energy. As this energy distribution flow is not the bottleneck of the generative task, we use the same preprocessing and architecture as in [31]. It consists of 6 MADE blocks [127, 128, 129] with rational quadratic splines (RQS) [130, 114]. We apply permutation layers between these MADE blocks. A logit transformation is used to preprocess the layer energies, and a log transformation is used for the incident energy.
@@ 3.3 Training
Calorimeter data are sparse, meaning that most input values are zero. This poses a problem for training flows. Since the data points live in a submanifold of the data space, the density and therefore the loss are not bounded. This problem is usually solved by adding noise. As in the original CaloFlow, we add uniform noise between zero and $1$ keV to all voxels. Only adding uniform noise would lead to discontinuities in the density at zero and one keV. To avoid that, we fill zero voxels with log Gaussian distributed values before adding the uniform noise.
We trained one instance of the networks for 200 epochs on GettingHigh and one instance for 800 epochs on CaloChallenge dataset 3. The different amount of epochs is necessary because of the different training statistics. We applied a one-cycle learning rate scheduler [135], decoupled L2 regularization [136], and L2 gradient clipping. ADAM [133] was used as optimizer. As data augmentation, we rotated the CaloChallenge dataset 3 showers by random numbers of angular bins in the $x-y$ plane. This is possible because the incident particles enter the calorimeter perpendicular to its surface and no magnetic field brakes the symmetry. Further details on architecture and training are provided in A.
@@ 3.4 Postprocessing
It is not guaranteed that the causal flows produces showers that match the layer energies generated by the energy distribution flow exactly. For that reason,[27, 31] rescaled showers per-layer to match the target. However, like the authors of [33], we observe that rescaling generated CaloChallenge showers results in an excess of low energy hits due to zero voxels getting shifted above the threshold. Therefore, instead of applying postprocessing for CaloChallenge, the showers generated by the causal flows are used directly. For GettingHigh we use the same energy rescaling as in [31]. It involves a way of identifying and masking empty voxels. See Appendix B.3 in [31] for a detailed description.
@@ 4.2 Classifier
Histograms and one-dimensional Wasserstein distances are useful to test how well one-dimensional distributions match, but they do not account for correlations. Instead, we use classifier tests to compare generative models based on higher-dimensional distributions [137, 27, 28, 33, 31].
@@ 5.1 Observables
The diagrams in Figure 9 show numerical evaluated one-dimensional Wasserstein distances between test data and generated data. Given are the Wasserstein distances for all observables described in section 4.1. To make the Wasserstein distance dimensionless, the distributions are standard scaled using the mean and standard deviation of the Geant4 distribution. We see a clear improvement compared to the BIB-AE in almost all observables. We also notice that for many observables, there are no differences between Geant4 and Flow-generated samples that can be resolved with the given amount of statistics.
@@ Acknowledgments
This research was supported in part by the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany. This project has received funding from the European Union’s Horizon 2020 Research and Innovation programme under Grant Agreement No 101004761. We acknowledge support by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and via the KISS consortium (05D23GU4, 13D22CH5) funded by the German Federal Ministry of Education and Research BMBF in the ErUM-Data action plan. CK would like to thank the Baden-Württemberg-Stiftung for financing through the program Internationale Spitzenforschung, project Uncertainties – Teaching AI its Limits (BWST_IF2020-010). DS is supported by the U.S. Department of Energy (DOE), Office of Science grant DOE-SC0010008.
@@ Appendix A Architecture and Training
In table 9, we give the total number of trainable parameters and the number of parameters involved in the generation process.
@@
model total generation
L2LFlows 163,269,320 163,269,320
BIB-AE 76,515,284 36,517,632
L2LFlows (CaloChallenge 3) 194,964,482 194,964,482
Table 9: Number of parameters of the models. middle: total number of parameters, right: number of parameters used for generation.
@@ Appendix B Preprocessing
$\text{logit}_{\alpha}(x)$ $=\log\left(\frac{x\times(1-2\alpha)+\alpha}{1-x\times(1-2\alpha)-\alpha}\right)$ (B.1)
@@
$E_{\text{max}}$ $=\max_{\text{training data}}\max_{i}E_{i}$ (B.2)
$x_{i}$ $=\text{logit}_{10^{-6}}\left(\frac{E_{i}}{E_{\text{max}}}\right)$ (B.3)
@@
$I_{\text{max},i}$ $=\max_{\text{training data}}\max_{jk}I_{ijk}$ (B.4)
$x_{ijk}$ $=\text{logit}_{10^{-2}}\left(1+\frac{\log\left(10^{-7}+\frac{I_{ijk}}{v_{\text{max},i}}\right)}{\log(10^{7})}\right)$ (B.5)
@@
$\text{condition}_{inc}$ $=\log\left(10^{-6}+\frac{E_{inc}}{1~{}\text{GeV}}\right)$ (B.6)
$\text{condition}_{i}$ $=\log\left(10^{-6}+\frac{E_{i}}{1~{}\text{GeV}}\right)$ (B.7)